Cletus has two vehicles: a 4x4 Ford pick-up truck that runs on diesel and a 1969 Dodge Charger that runs on gasoline. Suppose Cletus spends a fixed amount of money on automotive fuel each week. Cletus receives utility from buying diesel and gasoline because the fuels allow him to drive his vehicles.

Cletus currently purchases five gallons of gasoline per week and four gallons of diesel per week. Diesel costs $3 per gallon and gasoline costs $2 per gallon. At his current rate of consumption, Cletus spends all of his weekly fuel budget. His marginal utility (MU) from the fifth gallon of gas is 50 and his marginal utility from the fourth gallon of diesel is 60.
1.1. True or False: Cletus's current consumption levels of diesel and gasoline maximize his total utility.
1
Expert's answer
2017-10-17T15:58:07-0400
In order to the rule of maximizing utility:
MU1/P1= MU2/P2, where MU1 – marginal utility from the last gallon of gas, MU2 – marginal utility from the last gallon of diesel, P1 – price of all gallons of gas, P2 – price of all gallons of diesel.
50/(5*2)= 60/(4*3)=5
So, statement that Cletus's current consumption levels of diesel and gasoline maximize his total utility is true.